prove that no number of the type 4k+2 be a perfect square.
Answer by nabla(475)

Denote a^2=4k+2=2(2k+1). So 2|a^2 thus 2|a. Let a=2n. Then a^2=4n^2 which means 4|a^2.
However, we have a^2=4k+2. 4 does not divide 4k+2, hence 4k+2 can never be a perfect square.

In other words: if a is even, its square will always be divisible by 4.